OpenStudy (anonymous):

the ratio of three number is 3:4:5and the sum of the theire square is 1250.The sum of the three number is ?

OpenStudy (anonymous):

Let the numbers be 3x, 4x, 5x Thus according to the question \[(3x)^2 + (4x)^2 + (5x)^2 = 1250\] i.e. \[9x^2 + 16x^2 + 25x^2 = 1250\] i.e. \[50x^2 = 1250\] i.e. \[x^2 = \frac{1250}{50} = 25\] i.e. \[x =\pm 5\] Now if x=5 Then first number =3x= 15 second number =4x= 20 Third number =5x= 25 Sum of numbers = 15+ 20 + 25 = 60